# Chenjie Zhong    Zhipeng Li†,*    Shangzhi Xu Xiaohu Li

### v1

@@ 4.3 Experiments on Central Frequencies
Table 1 shows the result, where in each experiment the evaluated center frequencies for both ours and the pure VMD are sorted, and the max difference is computed as $\max\limits_{i}|f_{i}-F_{i}|$, where $f$ stands for our estimated frequencies and $F$ pure VMD’s. As shown in Table 1, in most cases, our estimation to the center frequencies has low difference with respect to the pure VMD. It can be found that, in Experiment 4, the error between estimated center frequencies of ours and pure VMD’s are larger than the average, however, as analysed in Experiment 4, on one side, the original VMD, no matter given IMFs number 2 or 3, fails to achieve reconstruction error as small as ours, on the other side, as shown in Fig.S5.13a(c) even when both started at IMFs number 3, the original VMD seems produce replicated modes at the Gaussian-alike component, with almost the same amplitude and shape, while ours provide relatively independent components when given our initial estimation for the center frequencies, which seems to indicates that our estimation in this case is more reliable.
@@ 6 Acknowledgement
SUPPLEMENTARY MATERIAL
1 Introduction of VMD
The key idea of VMD is to decompose an input signal into a series number of sub-signals via variational optimization method in spectral domain[5]. It is assumed the frequency of each mode is almost compact around a center pulsation and the associated analytic signal is computed by means of Hilbert Transform to get the unilateral frequency spectrum. Then each mode’s frequency spectrum was shifted into baseband by multiplying a frequency modulation factor with respect to the estimated center frequency, transforming each band-pass mode into low pass, and the bandwidth can be estimated by $H^{1}$ Gaussian smoothness. Afterwards, they decompose the original signal into each modes by minimizing the sum of the bandwidth of each mode with the constraints that the summation of all modes is the original signal. In short, their key variation optimization can be listed as following:
$\min\limits_{u_{k},\omega_{k}}\left\{\sum\limits_{k}\left\|\partial_{t}\left[\left(\delta(t)+\frac{\text{j}}{\pi t}\right)*u_{k}(t)\right]e^{-\text{j}\omega_{k}t}\right\|_{2}^{2}\right\}$
$\text{s.t.}\quad\sum\limits_{k}u_{k}=f$
so that each mode can be altered into an associated analytical signal by using Hilbert Transform and shifted into base-band for the following estimation of bandwidth using H1 Gaussian smoothness with the shifted demodulated components. To solve the problem, one can construct an augmented Lagrangian Function by introducing both quadratic penalty and Lagrangian Multipliers as follows:
$\mathcal{L}(u_{k},\omega_{k},\lambda)$ $=\alpha\sum_{k}\left\|\partial_{t}\left[\left(\delta(t)+\frac{j}{\pi t}\right)*u_{k}(t)\right]e^{-j\omega_{k}t}\right\|_{2}^{2}$
$u_{k}^{n+1}=\underset{u_{k}\in\mathbb{R}}{\arg\min}\left\{\alpha\left\|\partial_{t}\left[\left(\delta(t)+\frac{j}{\pi t}\right)*u_{k}(t)\right]e^{-j\omega_{k}t}\right\|_{2}^{2}+\left\|f-\sum u_{i}+\frac{\lambda}{2}\right\|_{2}^{2}\right\}.$
$\hat{u}_{k}^{n+1}=\underset{\hat{u}_{k},\,\hat{u}_{k}=\hat{u}_{k}^{T}}{\arg\min}\left\{\alpha\left\|j\omega\left[1+\text{sgn}(\omega+\omega_{k})\right]\hat{u}_{k}(\omega+\omega_{k})\right\|_{2}^{2}+\left\|\hat{f}-\sum\hat{u}_{i}+\frac{\lambda}{2}\right\|_{2}^{2}\right\}.$
$\hat{u}_{k}^{n+1}=\underset{u_{k},\,\hat{u}_{k}=\hat{u}_{k}^{T}}{\arg\min}\left\{\int_{0}^{\infty}4\alpha(\omega-\omega_{k})^{2}|\hat{u}_{k}(\omega)|^{2}d\omega+2\int\left(\hat{f}-\sum\hat{u}_{i}+\frac{\hat{\lambda}}{2}\right)^{2}d\omega\right\}$
which can be evaluated by vanishing the first variation of positive frequencies for both signal components and center frequencies
$\hat{u}_{k}^{n+1}=\left(\hat{f}-\sum_{i\neq k}\hat{u}_{i}+\frac{\hat{\lambda}}{2}\right)\frac{1}{1+2\alpha(\omega-\omega_{k})^{2}}$
$\omega_{k}^{n+1}=\frac{\displaystyle\int_{0}^{\infty}\omega\left|\hat{u}_{k}(\omega)\right|^{2}d\omega}{\displaystyle\int_{0}^{\infty}\left|\hat{u}_{k}(\omega)\right|^{2}d\omega}$
2 Finite Discrete Expression and Their Matrices
@@
(S2.5) $\boldsymbol{G}^{(1)}=\cfrac{1}{12h}\begin{bmatrix}1&-8&0&8&-1&0&\cdots&\cdots&\cdots&0\\ 0&1&-8&0&8&-1&0&\cdots&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&\cdots&\cdots&0&1&-8&0&8&-1&0\\ 0&0&\cdots&\cdots&0&1&-8&0&8&-1\\ 12h&0&0&\cdots&\cdots&\cdots&\cdots&0&0&0\\ 0&0&0&\cdots&\cdots&\cdots&\cdots&0&0&12h\\ -12&12&0&0&\cdots&\cdots&\cdots&0&0&0\\ 0&0&0&\cdots&\cdots&\cdots&0&0&-12&12\end{bmatrix}$
(S2.6) $\boldsymbol{G}^{(2)}=\cfrac{1}{12h^{2}}\begin{bmatrix}-1&16&-30&16&-1&0&\cdots&\cdots&\cdots&0\\ 0&-1&16&-30&16&-1&0&\cdots&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&\cdots&\cdots&0&-1&16&-30&16&-1&0\\ 0&0&\cdots&\cdots&0&-1&16&-30&16&-1\\ 12h^{2}&0&0&\cdots&\cdots&\cdots&\cdots&0&0&0\\ 0&0&0&\cdots&\cdots&\cdots&\cdots&0&0&12h^{2}\\ -12h&12h&0&0&\cdots&\cdots&\cdots&0&0&0\\ 0&0&0&\cdots&\cdots&\cdots&0&0&-12h&12h\end{bmatrix}$
(S2.7) $\boldsymbol{G}^{(3)}=\cfrac{1}{2h^{3}}\begin{bmatrix}-1&2&0&-2&1&0&\cdots&\cdots&\cdots&0\\ 0&-1&2&0&-2&1&0&\cdots&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&\cdots&\cdots&0&-1&2&0&-2&1&0\\ 0&0&\cdots&\cdots&0&-1&2&0&-2&1\\ 2h^{3}&0&0&\cdots&\cdots&\cdots&\cdots&0&0&0\\ 0&0&0&\cdots&\cdots&\cdots&\cdots&0&0&2h^{3}\\ -2h^{2}&2h^{2}&0&0&\cdots&\cdots&\cdots&0&0&0\\ 0&0&0&\cdots&\cdots&\cdots&0&0&-2h^{2}&2h^{2}\end{bmatrix}$
(S2.8) $\boldsymbol{G}^{(4)}=\cfrac{1}{h^{4}}\begin{bmatrix}1&-4&6&-4&1&0&\cdots&\cdots&\cdots&0\\ 0&1&-4&6&-4&1&0&\cdots&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&\cdots&\cdots&0&1&-4&6&-4&1&0\\ 0&0&\cdots&\cdots&0&1&-4&6&-4&1\\ h^{4}&0&0&\cdots&\cdots&\cdots&\cdots&0&0&0\\ 0&0&0&\cdots&\cdots&\cdots&\cdots&0&0&h^{4}\\ -h^{3}&h^{3}&0&0&\cdots&\cdots&\cdots&0&0&0\\ 0&0&0&\cdots&\cdots&\cdots&0&0&-h^{3}&h^{3}\end{bmatrix}$
where the last four lines in matrices correspond to the boundary condition and the conversion matrices $\boldsymbol{A}^{(n)}$ have the same structure.
3 About Component-Wise Multiplication
Our component-wise multiplication is not necessarily the Hadamard production of matrices, in that we add some special rules in dealing with column vector, to say in detail, $\forall\boldsymbol{A}_{m\times n},\boldsymbol{v}_{m\times 1}$, we have
(S3.1) $\boldsymbol{v}\odot\boldsymbol{A}=\boldsymbol{A}\odot\boldsymbol{v}=\boldsymbol{A}*\boldsymbol{V}_{m\times n}$
where * stands for standard Hadamard production and each column of $\boldsymbol{V}$ is just a copy of $\boldsymbol{v}$. In this way, we can transform (13) into
(S3.2) $((2\boldsymbol{A^{(2)}\alpha})\odot\boldsymbol{G^{(2)}})g+((4\boldsymbol{A^{(1)}\alpha})\odot\boldsymbol{G^{(3)}})g+((2\boldsymbol{\alpha})\odot\boldsymbol{G^{(4)}})g=\boldsymbol{\beta-\lambda+\mu}$
By applying this rule we have $(\boldsymbol{A}\boldsymbol{a})\odot(\boldsymbol{B}\boldsymbol{b})=((\boldsymbol{A}\boldsymbol{a})\odot\boldsymbol{B})\boldsymbol{b}$ so that we can treat (13) as common matrix equation and get the solution.
4 Supplementary Material of the Proof
4.1 Proof of Convexity
@@
Let $\mathcal{L}(g,\lambda,\mu,\alpha,\beta)=J_{1}(g(x))+J_{2}(g(x))$, where $J_{1}(g(x))=\displaystyle\int_{\Omega}\alpha(x)g^{\prime\prime 2}(x)\text{d}x$ and $J_{2}(g(x))=\displaystyle-\int_{\Omega}\beta(x)g(x)\text{d}x$. For $J_{1}(g(x))$, we have
(S4.1) $\lambda J_{1}(g_{1}(x))+(1-\lambda)J_{1}(g_{2}(x))-J_{1}[\lambda g_{1}(x)+(1-\lambda)g_{2}(x)]$
@@
and $J_{2}(g(x))$ is obviously a linear funtional, so we have
(S4.2) $J[\lambda g_{1}(x)+(1-\lambda)g_{2}(x)]\leq\lambda J[g_{1}(x)]+(1-\lambda)J[g_{2}(x)]$
@@
4.2 Proof of the Zero Duality Gap
@@
(S4.6) $\mathcal{G}(\lambda^{*}(x),\mu^{*}(x)=\inf\limits_{g(x)}\mathcal{L}(g(x),\lambda^{*}(x),\mu^{*}(x))\leq\inf\limits_{g(x)}\mathcal{O}(g(x))=\mathcal{O}(g^{*}(x))$
@@
hold for all $(p(x),q(x),o)$ in $\mathcal{S}$. Note that at least $\forall p(x),q(x)\geq 0$, $\mathcal{O}(g^{*}(x))\geq o$, so that the right end in this case is always non-negative. To summarize all above, we must have $\lambda(x)\geq 0$ and $\mu(x)\geq 0$, which indicates we can find a solution in dual space making the dual-gap vanish.

### v2

@@ 4.2 Experiments on Central Frequencies
Table 1 shows the result, where in each experiment the evaluated center frequencies for both ours and the pure VMD are sorted, and the max difference is computed as $\max\limits_{i}|f_{i}-F_{i}|$, where $f$ stands for our estimated frequencies and $F$ pure VMD’s. As shown in Table 1, in most cases, our estimation to the center frequencies has low difference with respect to the pure VMD. It can be found that, in Experiment 4, the error between estimated center frequencies of ours and pure VMD’s are larger than the average, however, as analysed in Experiment 4, on one side, the original VMD, no matter given IMFs number 2 or 3, fails to achieve reconstruction error as small as ours, on the other side, as shown in Fig.7c even when both started at IMFs number 3, the original VMD seems produce replicated modes at the Gaussian-alike component, with almost the same amplitude and shape, while ours provide relatively independent components when given our initial estimation for the center frequencies, which seems to indicates that our estimation in this case is more reliable.
Figure 9: The original spectrums for signal $y(t)=6t+\displaystyle\sum\limits_{i=1}^{10}(13-i)\cos[(20+10i)\pi t]$.
Figure 10: The decomposed spectrums for signal $y(t)=6t+\displaystyle\sum\limits_{i=1}^{10}(13-i)\cos[(20+10i)\pi t]$ by our method.
4.3 Experiments Compared with Variational Mode Decomposition
We noticed that, the Successive Variational Mode Decomposition(SVMD) method is another competitive method that can also automatically determine the number of IMFs for VMD[12] without any prior knowledge although it also lacks of rigorous proof on the convergence. This method retrieves the modes one after another by recursively finding the $L$th mode from the residual, assuming the first L1L-1 modes are found after the last turn, until the final residual can be ignored with respect to some criterion set by the algorithm. However, since the method is in essence recursive, the accuracy for the next mode to be evaluated is highly depended on the residual determined by previous decomposition in all former turns, which will inevitably introduce accumulated error during the propagation. We used official MATLAB implementation on SVMD [25] and observed that for narrow-banded signals wherein the center frequencies of different components are very close, the SVMD method may either introduce some redundant modes or drop some significant modes in their decomposition results, yet the phenomenon is hardly observed in our algorithm.
Figure 11: The decomposed spectrums for signal $y(t)=6t+\displaystyle\sum\limits_{i=1}^{10}(13-i)\cos[(20+10i)\pi t]$ by SVMD under different stopping criteria by considering noise(northwest), exact decomposition(northeast), Bayesian Estimation(southwest), power of the last mode(southeast).
Figure 12: The original spectrums for signal of the 102th record in MIT-BIH Arrhythmia Database.
In this section, we pick 2 typical signals for illustration. The first one is $y(t)=6t+\displaystyle\sum\limits_{i=1}^{10}(13-i)\cos[(20+10i)\pi t]$, which we have investigated in the experiments for finding the supporting baseline. For the completeness of our comparison, we investigated all stopping criterion implemented in the source code of SVMD. The method considering about eliminating noise(northwest) drops the base-frequency and a batch of high frequency, while method for extract reconstruction, method based on Beyesian Estimation, or method considering the power of last mode all result in some redundant modes with low energy on the decomposed spectrum (see colored spectrum at the bottom in each subplot), which are not the essential modes that should included in the decomposition result (Fig. 11). In contrast, our method(Fig. 10) accurately extracts all necessary modalities without any redundant ones.
Figure 13: The decomposed spectrums for the V2 lead of the 102th record in MIT-BIH Arrhythmia Database by our method.
Figure 14: The decomposed spectrums for signal with respect to the V2 lead of the 102th record in MIT-BIH Arrhythmia Database, and decomposed by [25] under different stopping criteria by considering noise(northwest), Bayesian Estimation(southwest), power of the last mode(southeast). Note that the figure for the exact decomposition is missed since the SVMD failed to converge in 10 seconds.
The second one is ECG signals from V2 lead of the 102th record in MIT-BIH Arrhythmia Database [23, 24]. The experiment result also shows similar phenomenon as stated in the former experiment, the method for exact decomposition fails to converge in more than 10 seconds so we cannot draw the spectrum of the corresponding figure (Fig. 14). Our method in this case can also succeed in retrieving all the effective modes(Fig. 13).
4.4 Experiment on Real Signals
Figure 15: The source signal (blue line) and the reconstructed signal (orange line) from the sum of IMFs obtained from our algorithm with respect to MLII lead from the 100th record.
Figure 16: The source signal (blue line) and the reconstructed signal (orange line) from the sum of IMFs obtained from our algorithm with respect to V4 lead from the 124th record.
In this section we investigate the performance of our method on real signals. As ECG signals are typical ones that combined with several narrow-banded components, we choose ECG signals as real signals for verification. Our ECG signals derives from MIT-BIH Arrhythmia Database [23, 24], and we choose randomly two signals for analysis, wherein the first is the MLII lead from the 100th record and the second is the V4 lead from the 124th record for our experiments respectively. Since each original signal has about 65000 samples which makes redundant for our experiment, we choose the first 2000 samples and do down-sampling at the ratio 1:10 for computation efficiency. For the limitation of the page we put the course of iteration and the decomposition result in Supplementary Materials(Fig. S3.16, Fig. S3.17 and Fig. S3.18, Fig. S3.19 for iteration and decomposition result with respect to the first and second chosen signals respectively), and instead we analyse the correlation coefficient(CC, the higher the better) and the ratio between the power ratio(PR) with respect to the residual and the original signal(the lower the better), and illustrate the gap between the reconstructed signal from the sum of the IMFs that acquired by our algorithm and the source signal.
For real signals composed of narrow-banded components, our algorithm can obtain effective components, since on one side the gap between the source signal(solid line in Fig. 15 and Fig. 16) and reconstructed signal(dashed line in Fig. 15 and Fig. 16) is typically small, meaning that one can always revive the source signal with small error by adding all the IMFs obtained from algorithm. On the other side, as shown in Table 2, the sum of IMF components obtained by our algorithm can achieve high correlation coefficient(both over 0.99) and low power ratio with respect to residual and the source signal(at the order of $10^{-4}$), which indicates our algorithm can automatically retrieve almost all the meaningful components to the source signal and seldom ignore the significance of components with small energy.
Table 2: The number of IMFs acquired from our algorithm, the CC(Correlation Coefficients between the reconstructed signal and source signal) and the PR(Power Ratio with respect to the residual and the source signal) result.
Experiment Number of IMFs $C$ PR
MLII lead from the 100th record 15 0.9989 4.2635e-04
V4 lead from the 124th record 14 0.9959 2.0626e-04
@@ 6 Acknowledgement
Appendix A Finite Discrete Expression and Their Matrices
@@
(S1.5) $\boldsymbol{A}^{(1)}=\boldsymbol{G}^{(1)}=\cfrac{1}{12}\begin{bmatrix}1&-8&0&8&-1&0&\cdots&\cdots&\cdots&0\\ 0&1&-8&0&8&-1&0&\cdots&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&\cdots&\cdots&0&1&-8&0&8&-1&0\\ 0&0&\cdots&\cdots&0&1&-8&0&8&-1\end{bmatrix}$
(S1.6) $\boldsymbol{A}^{(2)}=\boldsymbol{G}^{(2)}=\cfrac{1}{12}\begin{bmatrix}-1&16&-30&16&-1&0&\cdots&\cdots&\cdots&0\\ 0&-1&16&-30&16&-1&0&\cdots&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&\cdots&\cdots&0&-1&16&-30&16&-1&0\\ 0&0&\cdots&\cdots&0&-1&16&-30&16&-1\end{bmatrix}$
(S1.7) $\boldsymbol{A}^{(3)}=\boldsymbol{G}^{(3)}=\cfrac{1}{2}\begin{bmatrix}-1&2&0&-2&1&0&\cdots&\cdots&\cdots&0\\ 0&-1&2&0&-2&1&0&\cdots&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&\cdots&\cdots&0&-1&2&0&-2&1&0\\ 0&0&\cdots&\cdots&0&-1&2&0&-2&1\end{bmatrix}$
(S1.8) $\boldsymbol{A}^{(4)}=\boldsymbol{G}^{(4)}=\begin{bmatrix}1&-4&6&-4&1&0&\cdots&\cdots&\cdots&0\\ 0&1&-4&6&-4&1&0&\cdots&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&\cdots&\cdots&0&1&-4&6&-4&1&0\\ 0&0&\cdots&\cdots&0&1&-4&6&-4&1\end{bmatrix}$
and after evaluating $(2\boldsymbol{A}^{(2)}\boldsymbol{\alpha})\odot\boldsymbol{G}^{(2)}+(4\boldsymbol{A}^{(1)}\boldsymbol{\alpha})\odot\boldsymbol{G}^{(3)}+2\boldsymbol{\alpha}\odot\boldsymbol{G}^{(4)}$, we introduce another 4 lines depicting the boundary value condition, noted as $\boldsymbol{B}$:
(S1.9) $\boldsymbol{B}=\begin{bmatrix}1&0&0&\cdots&\cdots&0&0&0\\ 0&0&0&\cdots&\cdots&0&0&1\\ -1&1&0&\cdots&\cdots&0&0&0\\ 0&0&0&\cdots&\cdots&0&-1&1\end{bmatrix}$
and (19) is then transformed into
(S1.10) $\begin{bmatrix}(2\boldsymbol{A}^{(2)}\boldsymbol{\alpha})\odot\boldsymbol{G}^{(2)}+(4\boldsymbol{A}^{(1)}\boldsymbol{\alpha})\odot\boldsymbol{G}^{(3)}+2\boldsymbol{\alpha}\odot\boldsymbol{G}^{(4)}\\ \boldsymbol{B}\end{bmatrix}\boldsymbol{g}=\boldsymbol{c}$
in which
(S1.11) $\boldsymbol{\alpha}=\begin{bmatrix}\alpha_{0}\\ \alpha_{1}\\ \vdots\\ \alpha_{N-1}\\ \alpha_{N}\end{bmatrix},\boldsymbol{g}=\begin{bmatrix}g_{0}\\ g_{1}\\ \vdots\\ g_{N-1}\\ g_{N}\end{bmatrix}\text{and}\ \ \boldsymbol{c}=\begin{bmatrix}h^{4}c_{2}\\ h^{4}c_{3}\\ \vdots\\ h^{4}c_{N-3}\\ h^{4}c_{N-2}\\ c_{0}\\ c_{N}\\ c_{1}-c_{0}\\ c_{N}-c_{N-1}\end{bmatrix}$
and $c_{i}=\beta_{i}-\lambda_{i}+\mu_{i}$. In this way the condition number of the matrix to be inverted during evaluation is more gentle.
Appendix B Supplementary Material of the Proof
B.1 Proof of Convexity
Let $\mathcal{O}(g(x))=\mathcal{O}_{1}(g(x))+\mathcal{O}_{2}(g(x))$, where
@@ B.1 Proof of Convexity
$\mathcal{O}_{1}(g(x))=\displaystyle\int_{\Omega}\alpha(x)g^{\prime\prime 2}(x)\text{d}x$ and $\mathcal{O}_{2}(g(x))=\displaystyle-\int_{\Omega}\beta(x)g(x)\text{d}x$. For $\mathcal{O}_{1}(g(x))$, we have
(S2.1) $\lambda\mathcal{O}_{1}(g_{1}(x))+(1-\lambda)J_{1}(g_{2}(x))-\mathcal{O}_{1}[\lambda g_{1}(x)+(1-\lambda)g_{2}(x)]$
@@
and $\mathcal{O}_{2}(g(x))$ is obviously a linear funtional, so we have
(S2.2) $\mathcal{O}[\lambda g_{1}(x)+(1-\lambda)g_{2}(x)]\leq\lambda\mathcal{O}[g_{1}(x)]+(1-\lambda)\mathcal{O}[g_{2}(x)]$
@@
B.2 Concavity of Dual Functional $\mathcal{G}$
In this part we give a short proof to the concavity of $\mathcal{G}$. According to the definition of $\mathcal{G}$, we have
(S2.3) $\begin{split}&\ \ \ \ \mathcal{G}(\theta\lambda_{1}(x)+(1-\theta)\lambda_{2}(x),\theta\mu_{1}(x)+(1-\theta)\mu_{2}(x))\\ &=\inf\limits_{g(x)\in\mathcal{D}}\mathcal{L}[g(x),\theta\lambda_{1}(x)+(1-\theta)\lambda_{2}(x),\theta\mu_{1}(x)+(1-\theta)\mu_{2}(x)]\\ &=\inf\limits_{g(x)\in\mathcal{D}}\mathcal{L}[\theta g(x)+(1-\theta)g(x),\theta\lambda_{1}(x)+(1-\theta)\lambda_{2}(x),\theta\mu_{1}(x)+(1-\theta)\mu_{2}(x)]\\ &=\inf\limits_{g(x)\in\mathcal{D}}\{\theta\mathcal{L}[g(x),\lambda_{1}(x),\mu_{1}(x)]+(1-\theta)\mathcal{L}[(g(x),\lambda_{2}(x),\mu_{2}(x)]\}\\ &\geq\theta\inf\limits_{g(x)\in\mathcal{D}}\mathcal{L}[g(x),\lambda_{1}(x),\mu_{1}(x)]+(1-\theta)\inf\limits_{g(x)\in\mathcal{D}}\mathcal{L}[g(x),\lambda_{2}(x),\mu_{2}(x)]\\ &=\theta\mathcal{G}(\lambda_{1}(x),\mu_{1}(x))+(1-\theta)\mathcal{G}(\lambda_{2}(x),\mu_{2}(x))\end{split}$
which concludes the concavity of $\mathcal{G}$.
B.3 Proof of the Zero Duality Gap
@@
(S2.7) $\mathcal{G}(\lambda^{*}(x),\mu^{*}(x))=\inf\limits_{g(x)}\mathcal{L}(g(x),\lambda^{*}(x),\mu^{*}(x))\leq\inf\limits_{g(x)}\mathcal{O}(g(x))=\mathcal{O}(g^{*}(x))$
@@
hold for all $(p(x),q(x),o)$ in $\mathcal{S}$. Note that for $p(x),q(x)\geq 0$, there exist points in $\mathcal{S}$ that makes $o\leq\mathcal{O}(g^{*}(x))$, so there exists points that makes the right side of (S2.12) non-negative, this forces $\lambda(x)\geq 0$ and $\mu(x)\geq 0$, which indicates we can find a solution in dual space making the dual-gap vanish.